# The Ultraviolet Type Ia Supernova CubeSat (UVIa): Science Motivation & Mission Concept

### v1

@@ Abstract
The Ultraviolet (UV) Type Ia Supernova CubeSat (UVIa) is a CubeSat/SmallSat mission concept that stands to test critical space-borne UV technology for future missions like the Habitable Worlds Observatory (HWO) while elucidating long-standing questions about the explosion mechanisms of Type Ia supernovae (SNe Ia). UVIa will observe whether any SNe Ia emit excess UV light shortly after explosion to test progenitor/explosion models and provide follow-up over many days to characterize their UV and optical flux variations over time, assembling a comprehensive multi-band UV and optical low-redshift anchor sample for upcoming high-redshift SNe Ia surveys (e.g., Euclid, Vera Rubin Observatory, Nancy Roman Space Telescope). UVIa’s mission profile requires it to perform rapid and frequent visits to newly discovered SNe Ia, simultaneously observing each SNe Ia in two UV bands (FUV: 1500 – 1800 Å and NUV: 1800 – 2400 Å) and one optical band (-band: 3000 – 4200 Å). In this study, we describe the UVIa mission concept science motivation, mission design, and key technology development.
@@ 1 Introduction
The Ultraviolet Type Ia Supernova CubeSat (UVIa) is a Cubesat/Smallsat telescope concept dedicated to rapid and frequent UV/optical follow-up of Type Ia Supernovae (SNe Ia) in a compact, cost-effective package (Figure 1, left). UVIa is designed as a generalized transient astronomy response platform, to follow up supernova events caught with modern and upcoming ground-based all-sky surveyors (like the Zwicky Transient Factory[1] (ZTF), the Legacy Survey of Space and Time[2] (LSST) on the Rubin Observatory, BlackGEM, and Argus among others) less than a day after they are discovered. UVIa’s prime science sample is type Ia supernovae (SNe Ia) and will monitor them for up to 60 days after detection, providing the first statistical UV sampling of SNe Ia light curves out to 100 Mpc.
@@
UVIa contemporaneously observes in three photometric channels: FUV (1500 – 1800 Å), NUV (1800 – 2400 Å) and -band (3000 – 4200 Å). UVIa is sensitive to point sources down to an apparent magnitude of 21.5 AB in both UV bands (and 19 mag AB in -band) at a signal-to-noise ratio (S/N) of 5111Assuming a total integration time of 1 hour.. Recent advances in UV detector and mirror coating technologies allow UVIa to hit its projected performance in the UV (Figure 1, right). UV-sensitive delta-doped Complementary Metal–Oxide–Semiconductor (CMOS) Imaging Sensors (CIS) with integrated metal dielectric filter (MDF) coatings are projected to provide 40% quantum efficiency (QE) in UV bandpasses traditionally only able to achieve 12% [3]), as well as solar (optical) suppression (10-5). At the same time, emerging advances in multi-layer UV mirror coatings provide a promising avenue to achieve high throughput, narrow-band imaging filters in the FUV and NUV, further rejecting red light that makes accurate UV photometry challenging.
@@
All UV technologies have been called out as Tier 1 – 3 priorities in NASA’s Astrophysics Technology Gap Priority List and support critical advances in UV technology that serve a variety of UV instrument possibilities aboard the Habitable Worlds Observatory (HWO). UVIa is designed for a Sun Synchronous (SSO), Low Earth orbit (LEO), and its science operations are executed over the course of 12–18 months. The mission concept is optimized in a compact format, designed such that the payload could fit within a variety of commercially off-the-shelf (COTS) 16-U CubeSat or larger spacecraft (s/c) options.
@@ 2 Science Motivation
Type Ia supernovae (SNe Ia) have historically been used as cosmological distance indicators and led to the discovery of the accelerating expansion of the universe driven by dark energy [4, 5]. The longstanding cosmological relevance of SNe Ia has been borne out of the empirical fact that their light curve shapes correlate with their intrinsic luminosity, making them “standardizable” candles [6]. Despite this success of using SNe Ia as cosmological distance indicators, current surveys are already limited by systematics for SNe Ia cosmology[7, 8, 9]. This will only get worse as next-generation surveys like LSST on the Vera Rubin Observatory and the high-redshift surveys on the Nancy Grace Roman telescope begin, both of which are imminent. To achieve the full potential of these programs, we must build a calibration sample in the UV for high-redshift surveys (a sample that currently does not exist) and push for a better understanding of the progenitors of these cosmological explosions.
@@ 2.1 The Need for UV: Empirical UV Light Curves for Cosmology
Constraining UV light curves is key to improving our SN Ia distance measurements. Observationally, the largest scatter in SNe Ia colors (even after calibration) is in the UV [10]: HST NUV spectra of a handful of SNe Ia show that the UV has significant scatter, even when the optical is nearly indistinguishable [10]. This variation in the UV is not included in cosmological training samples due to the lack of data in the UV with the necessary photometric calibration (e.g. [11]).
@@
We choose a threshold of 20% of the Roman sample that will be redshifted into the currently uncalibrated UV band to estimate where a mission should include passband to overcome this lack of data. Rose et al. 2021[12] define a fiducial survey strategy for Roman that we then use to estimate the number of objects that will have restframe UV flux in their observations. They employ a two-tier, wide/deep, survey strategy. The bluest filter for the wide survey is while for the deep survey it is . Figure 2 shows the Roman band-passes overlaid with spectra of SN 2011fe at different redshifts.
@@
For the deep survey, 20% of the SNe Ia are at . The blue wavelength limit of -band, the bluest band for the Deep survey, is 927nm. This corresponds to a restframe of 309 nm which is restframe -band. For the highest redshift SNe, well-calibrated -band data is a necessity for SN Ia templates. For both surveys, 20% of the sample being unusable without UV observations is non-negligible: we need a nearby sample of SNe Ia for UV templates so that Roman can reach its full potential in constraining cosmological parameters like the dark energy equation of state parameter, .
@@ 2.2 UV Probes of the SN Ia Progenitor Problem
This companion shocking effect has not observed at the levels predicted for the classic single degenerate model, ruling out red giant companions for the majority of SN Ia progenitors[35, 15]. There are, however, some cases of an early UV excess similar to what was predicted by Kasen 2010[33]. SN 2017cbv is one of the best observed SN Ia with an early UV excess[36]. Rather than the excess we would have expected from a red giant, SN 2017cbv’s UV excess was more in line with a 7 main sequence star. Figure 3 illustrates the excess seen from a companion like that of SN 2017cbv. A similar excess with comparable size donor star was also found in SN 2019yvq[37]. This implies that there must be at least some SN Ia that arise from binaries that have a companion that is substantially larger than a WD. The binary companion for such a sub-Chandraskhar mass explosion is the variable which determines the mass of helium donated to the primary WD before explosion. A larger companion, like a He-sub dwarf or the stripped core of the companion star (after multiple rounds of mass transfer[38]) could donate a large amount of helium before detonating, producing a double detonation. This He-star companion may not be compact like a normal C/O WD and could produce a UV excess like those predicted for red giants or main sequence stars. Observations in the UV within days of explosion provides a window to measure the otherwise hidden donor star in SNe Ia.
One almost completely unexplored regime to constrain SN Ia progenitors is the FUV. Currently, FUV observations only exist for a single SN Ia (SN 2011fe) [13]. UVIa FUV observations will provide crucial constraints on present-day models of SNe Ia: the FUV is extremely sensitive to the outermost layers of the progenitor system. Many of these models converge in the optical by maximum light making them indistinguishable. FUV observations can break this degeneracy distinguishing between the proposed explosion mechanisms providing sorely needed constraints. Combined FUV/NUV color observations will provide unprecedented measurements of the color evolution of SNe Ia for the full population and will be essential to distinguish between the physical processes that produce UV excesses.
@@ 2.5 SN Ia Brightness Estimates
As no observations in the FUV exist for SN Ia at early time, we observations of SN 2011fe in the NUV[45], extrapolating the continuum into our passband. Due to faintness in the FUV, we consider SNe Ia at 80 Mpc which would still include objects. At 80 Mpc, our anticipated FUV brightness a week before maximum is 21.5 mag AB. At later times (e.g., peak brightness), there is evidence that the FUV flux relative to the NUV drops considerably [13], but the overall brightness of the SN increases, counteracting some of this dropoff. Due to this uncertainty, we use this image depth to conservatively meet our goal of 10 objects in the FUV, still successfully increasing the current sample by an order of magnitude.
@@ 3.1 Instrument Overview
The UVIa scientific instrument (Figure 6) consists of a 3 imaging telescopes that cover photometric bands between 1500 – 4500 Å (FUV: 1500 – 1800 Å, NUV: 1800 – 2400 Å, -band: 3000 – 4500 Å). A ray trace of the UVIa telescopes is shown in Figure 6. All UVIa telescopes are double-offset Cassegrain telescopes by design, driven to fit within a compact payload volume for a 12U CubeSat while ensuring quality on-axis spatial resolution with a fairly wide field of view (FOV)[51]. Each UV channel mirror is coated with a multi-layer (ML) thin film coating of alternating aluminum fluoride and lanthanum flouride (AlF3/LaF3), deposited via atomic layer deposition (ALD), where the number of fluoride layers and thicknesses determine each UV bandpass[52, 53, 54]. The -band channel mirrors baseline a standard aluminum (Al) with a thin magnesium fluoride (MgF2) over-coating and the SLOAN -band transmission filter222Should the Rubin/LSST -band filter be commercially available, UVIa will use that in place of the SLOAN filter. in the optical path in front of the optical camera. All channels feed to their own Teledyne-e2v CIS120-10-LN (“Capella”) CMOS sensor, featuring a 2k x 2k focal plane with 10m pixels and low intrinsic noise (7 /pix read noise and 0.05 /pix/s dark current at an operating temperature T-35∘C). The -band CMOS is back-illuminated with UV-enhancement, with projected efficiency based on performance reported for the Ozone Mapping and Profiler Suite (OMPS) mission[55]. Each FUV and NUV CMOS is back-illuminated and delta-doped[56, 57] with a custom integrated metal dielectric filter (MDF) made up of 7 layers of Al and AlF3; both UV CMOS follow processes developed of the Star-Planet Activity Research CubeSat (SPARCS) to optimize UV efficiency in each FUV and NUV bandpass while suppressing out-of-band light detection[58]. For further details on the optical design and technology associated with UVIa’s custom UV mirror coatings and detector filters, see Ref. 51.
@@ 3.2 Anticipated Performance
UVIa’s projected performance is based on the presented optical design and modeled/measured efficiencies of all UV and optical components integrated into the telescope design. The projected performance takes into account the in-band and out-of-band signal from the source (SN Ia) and noise contributions from the detector (read noise (RN), dark current noise (DC)), and in-band and out-of-band contributions from stray light (zodiacal and interstellar scattered light), sky background (Earth UV airglow features), and host galaxy contamination. Each noise contribution includes additional margin on conservative laboratory measurements (when available) and simulations of astronomical and stray light sources, based on empirical data from previous/current facilities. Figure 7 shows the projected breakdown of total counts from each source contributing to the signal pixel. Below, we describe how the expected count rates from each source are tabulated.
SNe Ia in-band signal: To date, only one Type Ia supernova, SN2011fe, has a complete, observed FUV-through-optical spectrum[13]. We use this one example as a “template” spectral energy distribution (SED) and scale the measured flux to UVIa’s desired minimum sensitivity (21.5 mag333All stated apparent magnitudes use the AB magnitude convention. through the UV and 19.0 mag in -band) to predict the in-band count rates for UVIa’s faintest target sampling. This constitutes the anticipated minimum in-band science signal count rate per pixel in each UVIa channel. Count rates in all bands as a function of time are shown in Figure 7. It is worth noting that the only FUV coverage of a SNe Ia to-date was observed about 2 weeks after the initial detection of the supernova, 3–4 days after peak -band brightness[13]. This is likely not representative of early-time UV detections UVIa would make, but it gives a conservative estimate of the UV fluxes and SED slopes.
@@
Even at these levels, we need to calibrate out the out-of-band contribution. To accomplish this, we will characterize the integrated system response in the laboratory, ahead of on-orbit science operations. Using this and the concurrent optical observations from the ground, we can model the out-of-band contribution and remove it. We will further refine our red leak model and monitor the red leak for any significant variations or changes to the telescope behaviors (e.g., due to UV coating or MDF instabilities or degradation on-orbit over time) with frequent standard star observations.
Detector noise: The CIS120-LN detector noise characteristics are assumed uniform over all pixels The CIS120-LN RN is reported to be 7 electrons/pixel (e/pix) by Teledyne-e2v and has been measured with a RN between 4 – 6 e/pix during laboratory testing with spaceflight-ready electronics[60]. Dark current (DC) is intrinsic to silicon-based detectors and has well-calibrated curves as a function of temperature[58]. At device temperatures achieved with passive cooling strategies, we expect a nominal DC rate of 0.01 e/pix/sec. This results in a total DC noise contribution less than the RN at exposure times 1000 seconds. By contrast, UVIa baselines an exposure time of 600 seconds per image. To ensure conservative estimates on performance, we assume an addition 30–50% margin on both RN and DN to calculate UVIa’s nominal sensitivity. This margin also accounts for increased noise from higher thermal conditions, camera electronics, and/or elevated noise from exposure to high radiation environments over time than anticipated.

### v2

# Ultraviolet Type Ia Supernova Mission (UVIa): Science Motivation & Mission Concept

@@ Abstract
The Ultraviolet (UV) Type Ia Supernova Mission (UVIa) is a CubeSat/SmallSat concept that stands to test critical space-borne UV technology for future missions like the Habitable Worlds Observatory (HWO) while elucidating long-standing questions about the explosion mechanisms of Type Ia supernovae (SNe Ia). UVIa will observe whether any SNe Ia emit excess UV light shortly after explosion to test progenitor/explosion models and provide follow-up over many days to characterize their UV and optical flux variations over time, assembling a comprehensive multi-band UV and optical low-redshift anchor sample for upcoming high-redshift SNe Ia surveys (e.g., Euclid, Vera Rubin Observatory, Nancy Roman Space Telescope). UVIa’s mission profile requires it to perform rapid and frequent visits to newly discovered SNe Ia, simultaneously observing each SNe Ia in two UV bands (FUV: 1500 – 1800 Å and NUV: 1800 – 2400 Å) and one optical band (-band: 3000 – 4200 Å). In this study, we describe the UVIa mission concept science motivation and basic mission design. The UVIa mission concept has been submitted to the CubeSats category of the NASA ROSES Astrophysics Research & Analysis (APRA) program ($10M cost cap) and NASA Astrophysics Pioneers program ($20M cost cap).
@@ 1 Introduction
The Ultraviolet Type Ia Supernova Mission (UVIa) is a Cubesat/Smallsat concept dedicated to rapid and frequent UV/optical follow-up of Type Ia Supernovae (SNe Ia) in a compact, cost-effective ($20M) package (Figure 1, left). UVIa is designed as a generalized transient astronomy response platform, to follow up supernova events caught with modern and upcoming ground-based all-sky surveyors (like the Zwicky Transient Factory[1] (ZTF), the Legacy Survey of Space and Time[2] (LSST) on the Rubin Observatory[3], BlackGEM[4], and Argus[5], among others) less than a day after they are discovered. UVIa’s prime science sample is SNe Ia and will monitor them for up to 60 days after detection, providing the first statistical UV sampling of SNe Ia light curves out to 100 Mpc. UVIa both compliments and provides a unique perspective on SNe Ia for cosmology and explosion physics within the landscape of current and future UV capabilities: UVIa is designed to observe multi-color UV-to-optical photometry contemporaneously (which UVEX and ULTRASAT lack), while providing high-cadence sampling of the SNe Ia over many days after the initial acquisition (which HST cannot do) with adequate red light suppression to provide unbiased UV photometry (which Swift cannot provide). UVIa compliments ULTRASAT’s transient discovery objective by offering a reactive, all-sky capability which focuses on a large fraction of SNe Ia within the local cosmological volume, while ULTRASAT adopts a narrow-deep approach, observing more distant objects with limited connection to nearby rungs of the cosmological distance ladder. UVIa would also compliment UVEX’s Community Target of Opportunity (ToO) program, providing much-needed constraints on UV light curves to take full advantage of the UV capabilities of the MIDEX mission.
@@
UVIa contemporaneously observes in three photometric channels: FUV (1500 – 1800 Å), NUV (1800 – 2400 Å) and -band (3000 – 4200 Å). With a total integration time of 1 hour, achieved by stacking 20 images with 300 second exposures in each channel, UVIa is sensitive to point sources down to an apparent magnitude of 21.5 AB in both UV bands (and 19 mag AB in -band) at a signal-to-noise ratio (S/N) of 5. Advances in UV detector[6, 7, 8, 9] and mirror coating[10, 11, 12] technologies allow UVIa to hit its projected performance in the UV (Figure 1, right). UV-sensitive delta-doped Complementary Metal–Oxide–Semiconductor (CMOS) Imaging Sensors (CIS) with integrated metal dielectric filter (MDF) coatings are projected to provide 40% quantum efficiency (QE) in UV bandpasses traditionally only able to achieve 12% [13]), as well as solar (optical) suppression (10-5)[14, 9]. At the same time, emerging advances in multi-layer UV mirror coatings provide a promising avenue to achieve high throughput, narrow-band imaging filters in the FUV and NUV, further rejecting red light that makes accurate UV photometry challenging[15].
@@
All UV technologies have been called out as Tier 1 – 3 priorities in NASA’s Astrophysics Technology Gap Priority List[16] and support critical advances in UV technology that help make possible sensitive UV instruments, including imagers and photometers, for future UV space-based observatories, including the Habitable Worlds Observatory (HWO). UVIa is designed for a Sun Synchronous (SSO), Low Earth orbit (LEO), and its science operations are executed over the course of 12–18 months. The mission concept is optimized in a compact format, designed such that the payload could fit within a variety of commercially off-the-shelf (COTS) 12U/16U CubeSat or larger spacecraft (s/c) options.
@@ 2 Science Motivation
Type Ia supernovae (SNe Ia) have historically been used as cosmological distance indicators and led to the discovery of the accelerating expansion of the universe driven by dark energy [17, 18]. The longstanding cosmological relevance of SNe Ia has been borne out of the empirical fact that their light curve shapes correlate with their intrinsic luminosity, making them “standardizable” candles [19]. Despite this success of using SNe Ia as cosmological distance indicators, current surveys are already limited by systematics for SNe Ia cosmology[20, 21, 22]: observing more SNe Ia will not improve cosmological constraints without improving our inference of the distances to SNe Ia. The “Hubble tension” – a 5.9 discrepancy between measured with supernovae ( km s-1 Mpc-1) [23], and the value predicted by the CMB ( km s-1 Mpc-1) [24], illustrates that there are either systematic errors in the distance measurements, or new physics is required. Moreover, recent indications (e.g., from DESI[25]) suggest time-varying dark energy whose cosmological signature makes SNe Ia too bright (by about 0.05 mag) at , yet too faint (by about 0.03 mag) at 1 – 2, relative to a cosmological constant model. Rubin and Roman can confirm or reject this model at high significance, but only if we can control for potential SN Ia evolution over cosmic time. Rest-frame UV observations, sensitive to astrophysical variations in SNe Ia, are the way to do this, but we currently lack the needed low-redshift sample. The large, high-redshift samples from next-generation surveys like LSST on the Vera Rubin Observatory and the Nancy Grace Roman telescope will not be able to resolve these cosmological “crises” without improved light curve modeling in the UV. UVIa will provide this training sample!
@@ 2.1 The Need for UV: Empirical UV Light Curves for Cosmology
Constraining UV light curves is key to improving our SN Ia distance measurements. Observationally, the largest scatter in SNe Ia colors (even after matching the stretch for the optical light curves) is in the UV [26]: HST NUV spectra of a handful of SNe Ia show that the UV has flux variations at the level, even when the optical is nearly indistinguishable [26]. This variation in the UV is not included in cosmological training samples due to the lack of data in the UV with the necessary photometric calibration (e.g. [27]).
@@
As an illustration, we choose a threshold of 20% of the Roman sample that will be redshifted into the currently uncalibrated UV band to estimate where a mission should include passband to overcome this lack of data. Rose et al. 2021[28] define a fiducial survey strategy for Roman that we then use to estimate the number of objects that will have restframe UV flux in their observations. They employ a two-tier, wide/deep, survey strategy. The bluest filter for the wide survey is while for the deep survey it is . Figure 2 shows the Roman band-passes overlaid with spectra of SN 2011fe at different redshifts.
@@
For the deep survey, 20% of the SNe Ia are at . The blue wavelength limit of -band, the bluest band for the Deep survey, is 927nm. This corresponds to a restframe of 309 nm which is restframe -band. For the highest redshift SNe, well-calibrated -band data is a necessity for SN Ia templates. For both surveys, 20% of the sample being unusable without UV observations is non-negligible: simple simulations of the recovered cosmological parameters shows a 50% increase in the precision on with the full sample compared to the sample excluding the highest redshift objects. We need a nearby sample of SNe Ia for UV templates so that Roman can reach its full potential in constraining cosmology.
@@ 2.2 UV Probes of the SN Ia Progenitor Problem
This companion shocking effect has not been observed at the levels predicted for the classic single degenerate model, ruling out red giant companions for the majority of SN Ia progenitors[51, 31]. There are, however, some cases of an early UV excess similar to what was predicted by Kasen 2010[49]. SN 2017cbv is one of the best observed SN Ia with an early UV excess[52]. Rather than the excess we would have expected from a red giant, SN 2017cbv’s UV excess was more in line with a 7 main sequence star. Figure 3 illustrates the excess seen from a companion like that of SN 2017cbv. A similar excess with comparable size donor star was also found in SN 2019yvq[53]. This implies that there must be at least some SN Ia that arise from binaries that have a companion that is substantially larger than a WD. The binary companion for such a sub-Chandraskhar mass explosion is the variable which determines the mass of helium donated to the primary WD before explosion. A larger companion, like a He-sub dwarf or the stripped core of the companion star (after multiple rounds of mass transfer[54]) could donate a large amount of helium before detonating, producing a double detonation. This He-star companion may not be compact like a normal C/O WD and could produce a UV excess like those predicted for red giants or main sequence stars. Observations in the UV within days of explosion provides a window to measure the otherwise hidden donor star in SNe Ia.
One almost completely unexplored regime to constrain SN Ia progenitors is the FUV. Currently, FUV observations only exist for a single SN Ia (SN 2011fe), taken with HST[29]. No other instrument with FUV coverage has had the ability to respond rapidly enough to observe SNe. UVIa FUV observations will provide crucial constraints on present-day models of SNe Ia: the FUV is extremely sensitive to the outermost layers of the progenitor system. Many of these models converge in the optical by maximum light making them indistinguishable. FUV observations can break this degeneracy distinguishing between the proposed explosion mechanisms providing sorely needed constraints. Combined FUV/NUV color observations will provide unprecedented measurements of the color evolution of SNe Ia for the full population and will be essential to distinguish between the physical processes that produce UV excesses.
@@ 2.3 Sample Size
Despite many years of effort, the dominant explosion mechanism is still hotly debated. With a sample size of 50, UVIa would provide statistically significant measurement of the fraction of the dominant UV production channel. Sub-dominant channels may only provide upper limits, but even that would move the field forward. For individual SNe Ia that match the expected properties for less common explosion mechanisms, UVIa would perform detailed studies on the specific objects providing additional constraints on those explosion models.
@@ 2.5 SN Ia Brightness Estimates
As no observations in the FUV exist for SN Ia at early time, we adopt observations of SN 2011fe in the NUV[61], extrapolating the continuum into our passband. Due to faintness in the FUV, we consider SNe Ia at 80 Mpc which would still include objects. At 80 Mpc, our anticipated FUV brightness a week before maximum is 21.5 mag AB. At later times (e.g., peak brightness), there is evidence that the FUV flux relative to the NUV drops considerably [29], but the overall brightness of the SN increases, counteracting some of this dropoff. Due to this uncertainty, we use this image depth to conservatively meet our goal of 10 objects in the FUV, still successfully increasing the current sample by an order of magnitude.
@@ 3.1 Instrument Overview
The UVIa scientific instrument (Figure 6) consists of 3 imaging telescopes that cover photometric bands between 1500 – 4500 Å (FUV: 1500 – 1800 Å, NUV: 1800 – 2400 Å, -band: 3000 – 4500 Å). A ray trace of the UVIa telescopes is shown in Figure 6. All UVIa telescopes are double-offset Cassegrain telescopes by design, driven to fit within a compact payload volume for a 12U CubeSat while ensuring quality on-axis spatial resolution with a fairly wide field of view (FOV)[67]. Each UV channel mirror is coated with a multi-layer (ML) thin film coating of alternating aluminum fluoride and lanthanum flouride (AlF3/LaF3), deposited via atomic layer deposition (ALD), where the number of fluoride layers and thicknesses determine each UV bandpass[68, 69, 11]. The -band channel mirrors baseline a standard aluminum (Al) with a thin magnesium fluoride (MgF2) over-coating and the SLOAN -band transmission filter. We note that, should the Rubin/LSST -band filter be commercially available, UVIa will use that in place of the SLOAN filter. in the optical path in front of the optical camera. All channels feed to their own Teledyne-e2v CIS120-10-LN (“Capella”) CMOS sensor, featuring a 2k x 2k focal plane with 10m pixels and low intrinsic noise (7 /pix read noise and 0.05 /pix/s dark current at an operating temperature T-35∘C). The -band CMOS is back-illuminated with UV-enhancement, with projected efficiency based on performance reported for the Ozone Mapping and Profiler Suite (OMPS) mission[70]. Each FUV and NUV CMOS is back-illuminated and delta-doped[71, 6] with a custom integrated metal dielectric filter (MDF) made up of 7 layers of Al and AlF3; both UV CMOS follow processes developed of the Star-Planet Activity Research CubeSat (SPARCS) to optimize UV efficiency in each FUV and NUV bandpass while suppressing out-of-band light detection[9]. For further details on the optical design, trade spaces considered, and technology associated with UVIa’s custom UV mirror coatings and detector filters, see Ref. 67.
@@ 3.2 Anticipated Performance
UVIa’s projected performance is based on the presented optical design and modeled/measured efficiencies of all UV and optical components integrated into the telescope design. The projected performance takes into account the in-band and out-of-band signal from the source (SN Ia) and noise contributions from the detector (read noise (RN), dark current noise (DC)), and in-band and out-of-band contributions from stray light (zodiacal and interstellar scattered light), sky background (Earth UV airglow features), and host galaxy contamination. Each noise contribution includes additional margin on conservative laboratory measurements (when available) and simulations of astronomical and stray light sources, based on empirical data from previous/current facilities. The projected counts as a function of source includes red leak contributions where applicable (e.g., out-of-band signal sources) and jitter effects from the s/c (taken into account as smearing of the target signal over multiple pixels; see Ref. 67 for more details and the optical error budget, which takes into account anticipated effects of s/c jitter). Figure 7 shows the projected breakdown of total counts from each source contributing to the signal pixel. Below, we describe how the expected count rates from each source are tabulated.
SNe Ia in-band signal: To date, only one Type Ia supernova, SN2011fe, has a complete, observed FUV-through-optical spectrum[29]. We use this one example as a “template” spectral energy distribution (SED) and scale the measured flux to UVIa’s desired minimum sensitivity (21.5 mag (AB) through the UV and 19.0 mag (AB) in -band) to predict the in-band count rates for UVIa’s faintest target sampling. This constitutes the anticipated minimum in-band science signal count rate per pixel in each UVIa channel. Total signal counts in all bands as a function of time are shown in Figure 7. It is worth noting that the only FUV coverage of a SNe Ia to-date was observed about 2 weeks after the initial detection of the supernova, 3–4 days after peak -band brightness[29]. This is likely not representative of early-time UV detections UVIa would make, but it gives a conservative estimate of the UV fluxes and SED slopes.
@@
Even at these levels, UVIa will calibrate out the out-of-band contribution. To accomplish this, we will characterize the integrated system response in the laboratory, ahead of on-orbit science operations. Using this and the concurrent optical observations from the ground, we can model the out-of-band contribution and remove it. We will further refine our red leak model and monitor the red leak for any significant variations or changes to the telescope behaviors (e.g., due to UV coating or MDF instabilities or degradation on-orbit over time) with frequent standard star observations.
Detector noise: The CIS120-LN detector noise characteristics are assumed uniform over all pixels The CIS120-LN RN is reported to be 7 electrons/pixel (e/pix) by Teledyne-e2v and has been measured with a RN between 4 – 6 e/pix during laboratory testing with spaceflight-ready electronics[73]. Dark current (DC) is intrinsic to silicon-based detectors and has well-calibrated curves as a function of temperature[9]. At device temperatures achieved with passive cooling strategies, we expect a nominal DC rate of 0.01 e/pix/sec. This results in a total DC noise contribution less than the RN at exposure times 1000 seconds. By contrast, UVIa baselines an exposure time of 300 seconds per image – the total integration time for all three channels to meet the sensitivity requirements involves the stacking of shorter exposure times to build up the required total integration time.. To ensure conservative estimates on performance, we assume an addition 30–50% margin on both RN and DN to calculate UVIa’s nominal sensitivity. This margin also accounts for increased noise from higher thermal conditions, camera electronics, and/or elevated noise from exposure to high radiation environments over time than anticipated.